# Formatting Instructions for ICLR 2024 Conference Submissions

### v3

@@ 3.3 Simulations with incomplete data
The results of each of these simulations are shown in Fig. 2b. We see that the predictive performance of models trained with the MSE loss degrades as the entropy of the fitness function in the epistatic domain increases, regardless of the type of nonlinearity that is applied to the latent fitness function. This is in contrast to the models trained with the BT loss, which often achieve nearly perfect estimation of the fitness function even when the entropy of the fitness function in the epistatic domain approaches its maximum possible value of $L\log 2$. This demonstrates the key result that the MSE loss is sensitive to the density of the epistatic representation resulting from global epistasis (as implied by the analogy to CS), while the BT loss is robust to these effects. We additionally find that these results are maintained when the degree of epistatic interactions in the latent fitness function is changed (Appendix D.1) and when Gaussian noise is added to the observed fitness functions (Appendix D.2
Next we tested how training set size effects the predictive performance of models trained with MSE and BT losses on a fitness function affected by global epistasis. In order to do so, we sampled a single $L=10,K=2$ fitness function from the NK model and applied the nonlinearity $g(f)=\exp(10\cdot f)$ to produce an observed fitness function. Then, for each of a range of training set sizes between 25 and 1000, we randomly sampled a training set and fit models with MSE and BT losses using the same models and procedure as in the previous simulations. We repeated this process for 200 training set replicates of each size, and calculated both the Spearman and Pearson correlations between the resulting model predictions and true observed fitness values for all sequences in the sequence space.
Fig 2c shows the mean correlation values across all 200 replicates of each training set size. There are two important takeaways from this plot. First, the BT loss achieves higher Spearman correlations than the MSE loss in all data regimes. This demonstrates the general effectiveness of this loss to estimate fitness functions affected by global epistasis. Next, we see that models trained with BT loss converge to a maximum Spearman correlation faster than models trained with MSE loss do to a maximum Pearson correlation, which demonstrates that the difference in predictive performance between models trained with MSE and BT losses is not simply due to a result of the evaluation metric being more tailored to one loss than the other. This result also reinforces our claim that fitness functions affected by global epistasis require more data to learn effectively with MSE loss, as would be predicted by CS scaling laws. The BT loss on the other hand, while not performant with the Pearson metric as expected by a ranking loss, seems to overcome this barrier and can be used to effectively estimate a fitness function from a small amount of data, despite the effects of global epistasis.
@@ 3.4 FLIP benchmark results
Figure 3: Comparison between MSE, Margin, and Bradley-Terry losses on FLIP benchmark tasks using the CNN baseline model. Panels correspond to tasks using the (a) GB1, (b) AAV, and (c) Thermostability datasets. Bar height and error bars indicate the mean and standard deviation, respectively, of test set Spearman correlation over 10 random initializations of the model. Asterisks indicate statistically significant improvement over another loss function (* p<0.05, ** p<0.01, *** p<0.001, **** p<0.0001).
@@
We trained the CNN model on each split using the standard MSE loss as well as the Margin and BT contrastive losses. The mean and standard deviation of Spearman correlations between the model predictions and test set labels over 10 random restarts are shown in Fig. 3. By default, the FLIP datasets contain portions of sequences that are never mutated in any of the data (e.g., only 4 positions are mutated in the GB1 data, but the splits contain the full GB1 sequence of length 56). We found that including these unmodified portions of the sequence often did not improve, and sometimes hurt, the predictive performance of the CNN models while requiring significantly increased computational complexity. Therefore most of our results are reported using inputs that contain only sequence positions that are mutated in at least one train or test data point. We found that including the unmodified portions of sequences improved the performance for the Low-vs-High and 3-vs-Rest GB1 well splits, as well as the 1-vs-rest AAV split and so these results are reported in Fig. 3; in these cases we found both models trained with MSE and contrastive losses had improved performance.
The results in Fig. 3 show that using contrastive losses (and particularly the BT loss) consistently results in improved predictive performance across a variety of practically relevant fitness prediction tasks (for more detail, see Appendix E.1 for the results in table form). Further, in no case does the BT loss result in worse performance than MSE. Indeed, it is shown in Otwinowski et al. [30] that a GB1 landscape closely associated with that in the FLIP benchmark is strongly affected by global epistasis. Further, many of the FLIP training sets are severely undersampled in the sense of CS scaling laws, which is the regime in which differences between MSE and contrastive losses are most apparent when global epistasis is present, as shown in Fig. 2.
Although Spearman correlation is commonly used to benchmark models of fitness functions, in practical protein engineering settings it also important to consider the ability of the model to classify the sequences with the highest fitness. To test this, we calculated the “top 10% recall“ for models trained with the BT and MSE losses on the FLIP benchmark data, which measures the ability of the models to correctly classify the 10% of test sequences with the highest fitness in the test set [19]. These results are shown in Appendix E.2, where it can be seen that the BT loss consistently outperforms the MSE loss.
@@ 4 Discussion
Additionally, our simulations largely do not consider how models trained with contrastive losses may be affected by the complex measurement noise commonly seen in experimental fitness assays based on sequencing counts [10]. Although our simulations mostly do not consider the effects of noise (except for the simple Gaussian noise added in Appendix D.2), our results on the FLIP benchmark demonstrate that contrastive losses can be robust to the effects of noise in practical scenarios. Further, we show in Appendix F: that the BT loss can be robust to noise in a potentially pathological scenario. A more complete analysis of the effects of noise on contrastive losses would complement these results.
Appendix Appendix A: Simulation details
@@ A.3 Definition of the NK model
The NK model is an extensively studied random field model of fitness functions introduced by Kauffman & Weinberger [24]. In order to sample a fitness function from the NK model, first one chooses values of the parameters $L$, $K$, and $q$, which correspond to the sequence length, maximum degree of epistatic interaction, and size of the sequence alphabet, respectively. Next, one samples a “neighborhood" $\mathcal{V}^{[j]}$ for each position $j$ in the sequence, which represents the $K$ positions that interact with position $j$. Concretely, for each position $j=1,...,L$, $\mathcal{V}^{[j]}$ is constructed by sampling $K$ values from the set of positions $\{1,...,L\}\setminus j$ uniformly at random, without replacement. Now let $\mathcal{S}^{(L,q)}$ be the set of all $q^{L}$ sequences of length $L$ and alphabet size $q$. Given the sampled neighborhoods $\{\mathcal{V}^{[j]}\}_{j=1}^{L}$, the NK model assigns a fitness to every sequence in $\mathcal{S}^{(L,q)}$ through the following two steps:
@@
Appendix Appendix B: Additional complete data results
@@ B.1 Additional nonlinearities and latent fitness function
B.2 Comparison to quantile transformation
In order to provide a baseline for the results shown in Figure 1c, we attempted to recover the latent fitness function using quantile transformations rather than using the Bradley-Terry loss. In particular, we applied a quantile transformation the observed fitness data y to produce a vector $\textbf{f}_{q}$ that approximately follows either a uniform or Gaussian distribution. In both cases, we used 10 quantiles for discretization. The results of these baseline tests are shown in Figure 5, below. By comparing to the results in Figure 1c, we can see that the quantile transformations are not as effective at recovering the latent fitness function as the Bradley-Terry loss.
Figure 5: Complete data recovery from a quantile transformation to a (a) uniform distribution (b) Gaussian distribution. Plot descriptions are as in Figure 1c.
Appendix Appendix C: Additional examples of uncertainty principle
@@ B.2 Comparison to quantile transformation
Appendix Appendix D: Additional incomplete data results
@@
D.1 Additional nonlinearity and latent fitness function parameters
@@
D.2 Adding noise to the observed data
@@
Appendix Appendix E: Additional results on FLIP benchmark
E.1 Results in table form
Here we reproduce the FLIP benchmark results shown in Figure 3 in table form, such that these results can be compared easily to other work that uses the FLIP benchmark.
Data set
Split MSE Loss Margin Bradley-Terry
GB1
1-vs-rest $0.133\pm 0.150$ $0.068\pm 0.093$ $0.091\pm 0.093$
2-vs-rest $0.564\pm 0.026$ $0.549\pm 0.041$ $0.607\pm 0.009$
3-vs-rest* $0.814\pm 0.049$ $0.869\pm 0.002$ $0.880\pm 0.003$
Low-vs-High* $0.499\pm 0.010$ $0.472\pm 0.020$ $0.567\pm 0.013$
Sampled $0.930\pm 0.002$ $0.944\pm 0.004$ $0.951\pm 0.002$
AAV
Mut-Des $0.751\pm 0.006$ $0.763\pm 0.005$ $0.757\pm 0.007$
Des-Mut $0.806\pm 0.006$ $0.806\pm 0.007$ $0.832\pm 0.002$
1-vs-rest* $0.335\pm 0.117$ $0.591\pm 0.042$ $0.485\pm 0.078$
2-vs-rest $0.748\pm 0.010$ $0.737\pm 0.009$ $0.798\pm 0.003$
7-vs-rest $0.732\pm 0.003$ $0.730\pm 0.003$ $0.742\pm 0.003$
Low-vs-High $0.401\pm 0.006$ $0.407\pm 0.007$ $0.410\pm 0.009$
Sampled $0.927\pm 0.001$ $0.922\pm 0.001$ $0.933\pm 0.000$
Thermostability
Mixed $0.349\pm 0.011$ $0.370\pm 0.009$ $0.453\pm 0.007$
Human $0.511\pm 0.016$ $0.526\pm 0.009$ $0.589\pm 0.002$
Human-Cell $0.490\pm 0.021$ $0.517\pm 0.012$ $0.570\pm 0.004$
Table 1: Comparison between MSE, Margin, and Bradley-Terry losses on FLIP benchmark tasks using the CNN baseline model. Each row represents a data set and split combination. Numerical columns indicate the mean and standard deviation of test set Spearman correlation over 10 random initializations of the model. Asterisks indicate that unmodified portions of sequences were used in training data.
E.2 Analysis of tail performance
Here we calculate the "top 10% recall" in the test set for the models trained with the BT and MSE losses on the FLIP benchmark data. The top 10% recall is defined as the fraction of test set sequences with fitness above the 90th percentile in the test set that are among the sequences with the highest 10% of model predictions. For examples, if there are 1000 sequences in the test set, then we determine how many of the 100 sequences with the highest fitness in the test set are in the set of sequences with the 100 largest model predictions.
Data set
Split MSE Loss Bradley-Terry
GB1
1-vs-rest 0.097 $\pm$ 0.030 0.138 $\pm$ 0.051
2-vs-rest 0.250 $\pm$ 0.030 0.282 $\pm$ 0.008
3-vs-rest* 0.539 $\pm$ 0.084 0.664 $\pm$ 0.014
Low-vs-High* 0.381 $\pm$ 0.028 0.443 $\pm$ 0.024
Sampled 0.823 $\pm$ 0.009 0.816 $\pm$ 0.010
AAV
Mut-Des $0.288\pm 0.004$ $\textbf{0.307}\pm\textbf{0.005}$
Des-Mut $0.318\pm 0.013$ $\textbf{0.387}\pm\textbf{0.008}$
1-vs-rest* $0.052\pm 0.053$ $\textbf{0.143}\pm\textbf{0.049}$
2-vs-rest $\textbf{0.490}\pm\textbf{0.011}$ $0.457\pm 0.010$
7-vs-rest $0.694\pm 0.006$ $0.695\pm 0.006$
Low-vs-High $\textbf{0.180}\pm\textbf{0.009}$ $0.170\pm 0.006$
Sampled $0.650\pm 0.005$ $0.652\pm 0.010$
Thermostability
Mixed $0.636\pm 0.013$ $0.616\pm 0.011$
Human $0.382\pm 0.022$ $\textbf{0.405}\pm\textbf{0.015}$
Human-Cell $0.316\pm 0.018$ $\textbf{0.355}\pm\textbf{0.012}$
Table 2: Tail performance comparison between MSE and Bradley-Terry losses on FLIP benchmark tasks using the CNN baseline model. Each row represents a data set and split combination. Numerical columns indicate the mean and standard deviation of top 10% recall on the test set over 10 random initializations of the model. Asterisks indicate that unmodified portions of sequences were used in training data. Bold values indicate significant improvement ($p<0.05$) over the other loss.
Appendix Appendix F: Noisy toy simulation
@@ E.2 Analysis of tail performance
In order to construct this scenario, we sampled an $L=10$, $K=2$ binary fitness function $f(\textbf{x}$ from the NK model, and then let $I(\textbf{x})=0$ when $f(\textbf{x})<\text{med}(f)$ and 1 otherwise, where $\text{med}(f)$ is the median NK fitness of all sequences. We then added Gaussian noise with $\sigma=0.1$ to produce the observed noisy data $y$. We consider $I(\textbf{x})$ to be the true binary label of a sequence and $y$ to be the noisy label of the sequence. The relationship between the NK fitness, true labels and noisy labels is shown in Fig. 10a, below. Next, we randomly split the data into training and test sets containing 100 and 924 sequences and noisy labels, respectively. We used the training set to fit two neural network models, one using the MSE loss and the other using the BT loss. The models and training procedure were the same as described in Appendix A:. The predictions of these models on test sequences, compared to noisy labels, are shown in 10b. We next constructed Receiver Operating Characteristic (ROC) curves that test the ability of each model to classify sequences according to their true labels $I(\textbf{x})$ (Fig. 10c). These curves show that while the MSE loss does slightly outperform the BT loss in this classification task, the BT loss still results in an effective classifier and does not exhibit a major pathology in this scenario.
@@
Appendix Appendix G: Derivation of fitness-epistasis uncertainty principle
@@
where $\otimes$ denotes the Kronecker product, and $\textbf{P}(q)$ is an orthonormal set of eigenvectors of the Graph Laplacian of the complete graph of size $q$. Based on Eq.  6, we can make the simplification to the calculation of the maximum value in $\bm{\Phi}$ that
@@
where I is the identity matrix and $\textbf{w}=\textbf{1}-\sqrt{q}\textbf{e}_{1}$, with 1 representing a vector with all elements equal to one, and $\textbf{e}_{1}$ is the vector equal to one in its first element and zero in all other elements. The values of each element in P(q) can be straightforwardly calculated using Eq.  8. In particular, we have:
@@
Putting these results together with Eq.  7, the fitness-epistasis uncertainty principle simplifies to

### v4

@@ 3.3 Simulations with incomplete data
The results of each of these simulations are shown in Fig. 2b. We see that the predictive performance of models trained with the MSE loss degrades as the entropy of the fitness function in the epistatic domain increases, regardless of the type of nonlinearity that is applied to the latent fitness function. This is in contrast to the models trained with the BT loss, which often achieve nearly perfect estimation of the fitness function even when the entropy of the fitness function in the epistatic domain approaches its maximum possible value of $L\log 2$. This demonstrates the key result that the MSE loss is sensitive to the density of the epistatic representation resulting from global epistasis (as implied by the analogy to CS), while the BT loss is robust to these effects. We additionally find that these results are maintained when the degree of epistatic interactions in the latent fitness function is changed (Appendix E.1) and when noise is added to the observed fitness functions (Appendix E.2).
Next, we tested how training set size effects the predictive performance of models trained with MSE and BT losses on a fitness function affected by global epistasis. In order to do so, we sampled a single $L=10,K=2$ fitness function from the NK model and applied the nonlinearity $g(f)=\exp(10\cdot f)$ to produce an observed fitness function. Then, for each of a range of training set sizes between 25 and 1000, we randomly sampled a training set and fit models with MSE and BT losses using the same models and procedure as in the previous simulations. We repeated this process for 200 training set replicates of each size, and calculated both the Spearman and Pearson correlations between the resulting model predictions and true observed fitness values for all sequences in the sequence space.
Fig. 2c shows the mean correlation values across all 200 replicates of each training set size. There are two important takeaways from this plot. First, the BT loss achieves higher Spearman correlations than the MSE loss in all data regimes. This demonstrates the general effectiveness of this loss to estimate fitness functions affected by global epistasis. Next, we see that models trained with BT loss converge to a maximum Spearman correlation faster than models trained with MSE loss do to a maximum Pearson correlation, which demonstrates that the difference in predictive performance between models trained with MSE and BT losses is not simply due to a result of the evaluation metric being more tailored to one loss than the other. This result also reinforces our claim that fitness functions affected by global epistasis require more data to learn effectively with MSE loss, as would be predicted by CS scaling laws. The BT loss on the other hand, while not performant with the Pearson metric as expected by a ranking loss, seems to overcome this barrier and can be used to estimate a fitness function from a small amount of data, despite the effects of global epistasis.
@@ 3.4 FLIP benchmark results
Spearman Top 10% Recall
Data set Split MSE Loss Bradley-Terry MSE Loss Bradley-Terry
GB1 1-vs-rest $0.133\pm 0.150$ $0.091\pm 0.093$ 0.097 $\pm$ 0.030 0.138 $\pm$ 0.051
2-vs-rest $0.564\pm 0.026$ $\textbf{0.607}\pm\textbf{0.009}$ 0.250 $\pm$ 0.030 0.282 $\pm$ 0.008
3-vs-rest* $0.814\pm 0.049$ $\textbf{0.880}\pm\textbf{0.003}$ 0.539 $\pm$ 0.084 0.664 $\pm$ 0.014
Low-vs-High* $0.499\pm 0.010$ $\textbf{0.567}\pm\textbf{0.013}$ 0.381 $\pm$ 0.028 0.443 $\pm$ 0.024
Sampled $0.930\pm 0.002$ $\textbf{0.951}\pm\textbf{0.002}$ 0.823 $\pm$ 0.009 0.816 $\pm$ 0.010
AAV Mut-Des $0.751\pm 0.006$ $0.757\pm 0.007$ $0.288\pm 0.004$ $\textbf{0.307}\pm\textbf{0.005}$
Des-Mut $0.806\pm 0.006$ $\textbf{0.832}\pm\textbf{0.002}$ $0.318\pm 0.013$ $\textbf{0.387}\pm\textbf{0.008}$
1-vs-rest* $0.335\pm 0.117$ $\textbf{0.485}\pm\textbf{0.078}$ $0.052\pm 0.053$ $\textbf{0.143}\pm\textbf{0.049}$
2-vs-rest $0.748\pm 0.010$ $\textbf{0.798}\pm\textbf{0.003 }$ $\textbf{0.490}\pm\textbf{0.011}$ $0.457\pm 0.010$
7-vs-rest $0.732\pm 0.003$ $\textbf{0.742}\pm\textbf{0.003}$ $0.694\pm 0.006$ $0.695\pm 0.006$
Low-vs-High $0.401\pm 0.006$ $\textbf{0.410}\pm\textbf{0.009}$ $\textbf{0.180}\pm\textbf{0.009}$ $0.170\pm 0.006$
Sampled $0.927\pm 0.001$ $\textbf{0.933}\pm\textbf{0.000}$ $0.650\pm 0.005$ $0.652\pm 0.010$
Thermostability Mixed $0.349\pm 0.011$ $\textbf{0.453}\pm\textbf{0.007}$ $0.636\pm 0.013$ $0.616\pm 0.011$
Human $0.511\pm 0.016$ $\textbf{0.589}\pm\textbf{0.002}$ $0.382\pm 0.022$ $\textbf{0.405}\pm\textbf{0.015}$
Human-Cell $0.490\pm 0.021$ $\textbf{0.570}\pm\textbf{0.004}$ $0.316\pm 0.018$ $\textbf{0.355}\pm\textbf{0.012}$
Table 1: Comparison between MSE and Bradley-Terry losses on FLIP benchmark tasks using the CNN baseline model. Each row represents a data set and split combination. Numerical columns indicate the mean and standard deviation of test set metrics over 10 random initializations of the model. Asterisks indicate that unmodified portions of sequences were used in training data. Bold values indicate that a loss has significantly improved performance over all other tested losses ($p<0.05$).
@@
We trained the CNN model on each split using the standard MSE loss and BT contrastive losses. The mean and standard deviation of Spearman correlations between the model predictions and test set labels over 10 random restarts are shown in Table 1, third and fourth columns. By default, the FLIP datasets contain portions of sequences that are never mutated in any of the data (e.g., only 4 positions are mutated in the GB1 data, but the splits contain the full GB1 sequence of length 56). We found that including these unmodified portions of the sequence often did not improve, and sometimes hurt, the predictive performance of the CNN models while requiring significantly increased computational complexity. Therefore most of our results are reported using inputs that contain only sequence positions that are mutated in at least one train or test data point. We found that including the unmodified portions of sequences improved the performance for the Low-vs-High and 3-vs-Rest GB1 well splits, as well as the 1-vs-rest AAV split and so these results are reported in Table 1; in these cases we found both models trained with MSE and contrastive losses had improved performance.
Although Spearman correlation is commonly used to benchmark models of fitness functions, in practical protein engineering settings it also important to consider the ability of the model to classify the sequences with the highest fitness. To test this, we calculated the “top 10% recall“ for models trained with the BT and MSE losses on the FLIP benchmark data, which measures the ability of the models to correctly classify the 10% of test sequences with the highest fitness in the test set [18]. These results are shown in the fifth and sixth columns of Table 1.
The results in Table1 show that using contrastive losses (and particularly the BT loss) consistently results in improved predictive performance across a variety of practically relevant fitness prediction tasks. Further, in no case does the BT loss result in worse performance than MSE. The reasons for this result may be manifold; however, we hypothesize that it is partially a result of sparse latent fitness functions being corrupted by global epistasis Indeed, it is shown in Otwinowski et al. [31] that a GB1 landscape closely associated with that in the FLIP benchmark is strongly affected by global epistasis. Further, many of the FLIP training sets are severely undersampled in the sense of CS scaling laws, which is the regime in which differences between MSE and contrastive losses are most apparent when global epistasis is present, as shown in Fig. 2. In order to ensure that our results translate to additional types of fitness data outside of the FLIP benchmark, we also compared the BT and MSE losses on two of the datasets from the ProteinGym benchmark [28]; these corroborating results are shown in Appendix G.
@@ 4 Discussion
Additionally, our simulations largely do not consider how models trained with contrastive losses may be affected by the complex measurement noise commonly seen in experimental fitness assays based on sequencing counts [10]. Although our simulations mostly do not consider the effects of noise (except for the simple Gaussian noise added in Appendix E.2), our results on the FLIP benchmark demonstrate that contrastive losses can be robust to the effects of noise in practical scenarios. Further, we show in Appendix F that the BT loss can be robust to noise in a potentially pathological scenario. A more complete analysis of the effects of noise on contrastive losses would complement these results.
Appendix Appendix A Simulation details
@@ A.3 Definition of the NK model
The NK model is an extensively studied random field model of fitness functions introduced by Kauffman and Weinberger [24]. In order to sample a fitness function from the NK model, first one chooses values of the parameters $L$, $K$, and $q$, which correspond to the sequence length, maximum degree of epistatic interaction, and size of the sequence alphabet, respectively. Next, one samples a “neighborhood” $\mathcal{V}^{[j]}$ for each position $j$ in the sequence, which represents the $K$ positions that interact with position $j$. Concretely, for each position $j=1,...,L$, $\mathcal{V}^{[j]}$ is constructed by sampling $K$ values from the set of positions $\{1,...,L\}\setminus j$ uniformly at random, without replacement. Now let $\mathcal{S}^{(L,q)}$ be the set of all $q^{L}$ sequences of length $L$ and alphabet size $q$. Given the sampled neighborhoods $\{\mathcal{V}^{[j]}\}_{j=1}^{L}$, the NK model assigns a fitness to every sequence in $\mathcal{S}^{(L,q)}$ through the following two steps:
@@
Appendix Appendix B Additional complete data results
@@ B.1 Additional nonlinearities and latent fitness function
Appendix Appendix C Characterizing nonlinearities that decrease entropy
Here we provide a sufficient condition on a nonlinearity for the nonlinearity to reduce the entropy of a fitness function in the fitness domain. In other words, we characterize a class of $g$ such that $H(g(\textbf{f}))\leq H(\textbf{f})$. First, we define probability vectors p and q such that $p_{i}\coloneqq\dfrac{f_{i}^{2}}{||\textbf{f}||^{2}}$ and $q_{i}\coloneqq\dfrac{g(f_{i})^{2}}{||g(\textbf{f})||^{2}}$ for all $i=1,...,N$.
We additionally define the vector w such that $w_{i}\coloneqq\dfrac{q_{i}}{p_{i}}$ and the probability vector $\textbf{r}(\alpha)$ where $r_{i}(\alpha)=\dfrac{p_{i}w_{i}^{\alpha}}{\sum_{j=1}^{N}p_{i}w_{i}^{\alpha}}$ for all $i=1,...,N$. Note that $\textbf{r}(0)=\textbf{p}$ and $\textbf{r}(1)=\textbf{q}$.
Finally, let $H_{s}(\textbf{p})=-\sum_{i=1}^{N}p_{i}\log p_{i}$ be the Shannon entropy of a probability vector p.
A note on notation: below we are primarily concerned with distributions over discrete outcomes indexed by integers. Given a probability vector p representing such a distribution and a vector x representing values associated with each outcome, we denote the expectation of values of x over the distribution represented by p as $E_{p}[x]=\sum_{i}p_{i}x_{i}$. Similarly, we write covariances as $\textnormal{Cov}_{p}(x,y)=E_{p}[xy]-E_{p}[x]E_{p}[y]$.
Theorem 1.
Assume without loss of generality that p is sorted such that $p_{i}\geq p_{i+1}$ for $i=1,...,N-1$. Additionally assume that $q_{i}=0$ if $p_{i}=0$ for $i$. If $w_{i}\geq w_{i+1}$ for all $i=1,...,N-1$, then $H(g(\textbf{f}))\leq H(\textbf{f})$.
To prove Theorem 1, we first provide a number of lemmas.
Lemma 1.
The derivative of the Shannon entropy of $\textbf{r}(\alpha)$ with respect to $\alpha$ is equal to the negative covariance between $\log\textbf{r}(\alpha)$ and $\log\textbf{w}$. Specifically, the derivative is given by
$\dfrac{\mathrm{d}H_{s}(\textbf{r}(\alpha))}{\mathrm{d}\alpha}=-\textnormal{Cov}_{r(\alpha)}(\log r(\alpha),\log w).$
Lemma 2.
Given a probability vector p of length $N$ and any two vectors x and y of length $N$, we have
$\textnormal{Cov}_{p}(x,y)=\sum_{j>i}p_{i}p_{j}(x_{i}-x_{j})(y_{i}-y_{j}).$ (5)
Corollary 1.
If two vectors $x$ and $y$ of length $N$ are sorted such that $x_{i}\geq x_{i+1}$ and $y_{i}\geq y_{i+1}$ for $i=1,...,N-1$, then $\textnormal{Cov}_{p}(x,y)\geq 0$ for any probability vector p.
Proof of Theorem 1.
First we note that by the definition of entropy given in the main text, we have $H(\textbf{f})=H_{s}(\textbf{p})$ and $H(g(\textbf{f}))=H_{s}(\textbf{q})$. Therefore we need only prove that $H_{s}(\textbf{q})\leq H_{s}(\textbf{p})$.
By Lemma 1, we have that
$H_{s}(\textbf{q})-H_{s}(\textbf{p})=-\int_{0}^{1}\textnormal{Cov}_{r(\alpha)}(\log r(\alpha),\log w)\mathrm{d}\alpha.$ (6)
Both p and w are sorted by assumption and therefore it is clearly true that $\textbf{r}(\alpha)$ is also sorted such that $r_{i}(\alpha)\geq r_{i+1}(\alpha)$ for $i=1,...,N-1$ and all $\alpha\in[0,1]$. Since the logarithm is a monotonic function, both $\log\textbf{r}(\alpha)$ and $\log\textbf{w}$ are similarly sorted. Therefore, by Corollary 1, we have that the integrand in Eq.  6 is positive for all $\alpha\in[0,1]$. Therefore, the integral must be positive and $H_{s}(\textbf{q})-H_{s}(\textbf{p})<0$.
Proof of Lemma 1.
$\dfrac{\mathrm{d}H_{s}(\textbf{r}(\alpha))}{\mathrm{d}\alpha}$ $=-\sum_{i=1}^{N}(1+\log r_{i}(\alpha))\dfrac{\mathrm{d}r_{i}(\alpha)}{\mathrm{d}\alpha}$
$=-\sum_{i=1}^{N}\left(1+\log r_{i}(\alpha)\right)\dfrac{p_{i}w_{i}^{\alpha}}{\sum_{j=1}^{N}p_{i}w_{i}^{\alpha}}\left(\log w_{i}-\dfrac{\sum_{k=1}^{N}p_{k}w_{k}^{\alpha}\log w_{k}}{\sum_{j=1}^{N}p_{i}w_{i}^{\alpha}}\right)$
$=-\sum_{i=1}^{N}r_{i}(\alpha)(1+\log r_{i}(\alpha))\left(\log w_{i}-\sum_{k}r_{k}(\alpha)\log w_{k}\right)$
$=-\sum_{i=1}^{N}r_{i}(\alpha)\left(\log w_{i}+\log r_{i}(\alpha)\log w_{i}-E_{r(\alpha)}[\log w]-\log r_{i}(\alpha)E_{r(\alpha)}[\log w]\right)$
$=-E_{r(\alpha)}[\log w]-E_{r(\alpha)}[\log r(\alpha)\log w]+E_{r(\alpha)}[\log w]+E_{r(\alpha)}[\log r(\alpha)]E_{r(\alpha)}[\log w]$
$=-E_{r(\alpha)}[\log r(\alpha)\log w]+E_{r(\alpha)}[\log r(\alpha)]E_{r(\alpha)}[\log w]$
$=-\text{Cov}_{r(\alpha)}(\log r(\alpha),\log w)$
Proof of Lemma 2.
$\sum_{i=1}^{N}\sum_{j=i}^{N}p_{i}p_{j}(x_{i}-x_{j})(y_{i}-y_{j})=\dfrac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}p_{i}p_{j}(x_{i}-x_{j})(y_{i}-y_{j})$
$=\dfrac{1}{2}\sum_{i=1}^{N}p_{i}x_{i}y_{i}\sum_{j=1}^{N}p_{j}-\dfrac{1}{2}\sum_{i=1}^{N}p_{i}x_{i}\sum_{j=1}^{N}p_{j}y_{j}-\dfrac{1}{2}\sum_{i=1}^{N}p_{i}y_{i}\sum_{j=1}^{N}p_{j}x_{j}+\dfrac{1}{2}\sum_{j=1}^{N}p_{j}x_{j}y_{j}\sum_{i=1}^{N}p_{i}$
$=E_{p}[xy]-E_{p}[x]E_{p}[y]$
$=\textnormal{Cov}_{p}(x,y)$
where the first line follows from the summand equaling zero when $i=j$. ∎
Proof of Corollary 1.
This is a straightforward consequence of Lemma 2. For every term in the sum of Eq.  5, we have that $j>i$, and therefore $x_{i}-x_{j}\geq 0$ and $y_{i}-y_{j}\geq 0$ by assumption. Further p is a probability vector and thus $p_{i}>0$ for all $i$. Therefore every element of the sum is positive and the covariance is positive.
Notably, the condition of Theorem 1 is not necessary in order decrease entropy in the fitness domain. Indeed, many of the nonlinearities tested herein do not satisfy this condition. Eq.  6 is an illuminating relationship that may provide further insight into additional classes of nonlinearities that decrease entropy.
Appendix Appendix D Additional examples of uncertainty principle
@@
Appendix Appendix E Additional incomplete data results
@@
E.1 Additional nonlinearity and latent fitness function parameters
@@
E.2 Adding noise to the observed data
@@
Appendix Appendix F Noisy toy simulation
@@
In order to construct this scenario, we sampled an $L=10$, $K=2$ binary fitness function $f(\textbf{x}$ from the NK model, and then let $I(\textbf{x})=0$ when $f(\textbf{x})<\text{med}(f)$ and 1 otherwise, where $\text{med}(f)$ is the median NK fitness of all sequences. We then added Gaussian noise with $\sigma=0.1$ to produce the observed noisy data $y$. We consider $I(\textbf{x})$ to be the true binary label of a sequence and $y$ to be the noisy label of the sequence. The relationship between the NK fitness, true labels and noisy labels is shown in Fig. 8a, below. Next, we randomly split the data into training and test sets containing 100 and 924 sequences and noisy labels, respectively. We used the training set to fit two neural network models, one using the MSE loss and the other using the BT loss. The models and training procedure were the same as described in Appendix A. The predictions of these models on test sequences, compared to noisy labels, are shown in 8b. We next constructed Receiver Operating Characteristic (ROC) curves that test the ability of each model to classify sequences according to their true labels $I(\textbf{x})$ (Fig. 8c). These curves show that while the MSE loss does slightly outperform the BT loss in this classification task, the BT loss still results in an effective classifier and does not exhibit a major pathology in this scenario.
@@
Appendix Appendix G Results on ProteinGym benchmark datasets
We compared the MSE vs Bradley-Terry loss on the CAPSD_AAV2S_Sinai_2021 and GFP_AEQVI_Sarkisyan_2016 datasets from the ProteinGym benchmark [28] using the same protocol as described in 3.4. These are two of the most relevant datasets in ProteinGym because they contain a large number of multi-mutants, whereas many other datasets contain only single and double mutations. The results for 10 replicates of uniform 80/20 train test splits are shown below:
Spearman Top 10% Recall
Dataset MSE Loss Bradley-Terry MSE Loss Bradley-Terry
CAPSD_AAV2S_Sinai_2021 $0.912\pm 0.003$ $\textbf{0.920}\pm\textbf{0.003}$ 0.915 $\pm$ 0.001 0.915 $\pm$ 0.001
GFP_AEQVI_Sarkisyan_2016 $0.867\pm 0.001$ $\textbf{0.873}\pm\textbf{0.001}$ 0.938 $\pm$ 0.003 0.945 $\pm$ 0.003
Table 2: Comparison between MSE and Bradley-Terry losses on ProteinGym benchmark tasks using the CNN baseline model. Each row represents a data set and split combination. Numerical columns indicate the mean and standard deviation of test set metrics over 10 random initializations of the model. Asterisks indicate that unmodified portions of sequences were used in training data. Bold values indicate that a loss has significantly improved performance over all other tested losses ($p<0.05$).
Appendix Appendix H Derivation of fitness-epistasis uncertainty principle
@@
where $\otimes$ denotes the Kronecker product, and $\textbf{P}(q)$ is an orthonormal set of eigenvectors of the Graph Laplacian of the complete graph of size $q$. Based on Eq.  8, we can make the simplification to the calculation of the maximum value in $\bm{\Phi}$ that
@@
where I is the identity matrix and $\textbf{w}=\textbf{1}-\sqrt{q}\textbf{e}_{1}$, with 1 representing a vector with all elements equal to one, and $\textbf{e}_{1}$ is the vector equal to one in its first element and zero in all other elements. The values of each element in P(q) can be straightforwardly calculated using Eq.  10. In particular, we have:
@@
Putting these results together with Eq.  9, the fitness-epistasis uncertainty principle simplifies to